# Quantum Gravity, Hydrodynamics and Emergent Cosmology: A Collection of Perspectives

### v1

@@ 3.1 Effective dynamics for the spatial volume of 3-dimensional CDT
One feature of CDT is the existence of a phase characterized by the emergence of a macroscopic universe, suggestive of a semiclassical behavior at large scales. A spacetime condensation phenomenon underlies the emergence of such a semiclassical universe, where the time extension of the condensate, or droplet, is strictly smaller than the total time of the quantum universe. In Benedetti:2014dra ; Benedetti:2016rwo ; Benedetti:2022ots we introduced an effective model for the dynamics of the spatial volume in $2+1$ dimensions, which can be obtained as a minisuperspace reduction of Hořava-Lifshitz gravity, and we have shown that a study of the minima of the action (supported also by Monte-Carlo simulations) leads to a phase diagram with a droplet phase similar to that of CDT, which instead cannot be reproduced with a model based on a minisuperspace reduction of Einstein’s gravity theory. Therefore, in view also of previous results  Horava:2009if ; Benedetti:2009ge ; Ambjorn:2010hu ; Budd:2011zm ; Ambjorn:2013joa , we conclude that there is strong evidence that the effective field theory associated to the continuum limit of CDT in $1+1$ and $2+1$ dimensions is a foliation-preserving gravity theory, that is, a theory in the same general class to which Hořava-Lifshitz gravity belongs (see Benedetti:2022ots for a detailed argument).
@@ 3.3 Three Ways to Mean-Field Group Field Theory
Mean-field theory is applicable if ${d_{\textrm{eff}}}>{d_{\textrm{crit}}}$. This works even for compact group if there are sufficiently many matter degrees of freedom adding to ${d_{\textrm{eff}}}$. On groups with hyperbolic geometry, as natural when there is some Lorentzian structure in a TGFT model as in Jercher:2021bie ; Jercher:2022mky , mean-field theory applies generically as ${d_{\textrm{eff}}}(k)\to\infty$ excedes the critical dimension in any case Marchetti:2022igl ; Marchetti:2022nrf . This lays the theoretical foundations for applications of TGFT to phenomenology based on condensate states in a mean-field regime, in particular to TGFT condensate cosmology Gielen:2013kla ; Gielen:2013naa ; deCesare:2016rsf ; Pithis:2016wzf ; Pithis:2016cxg ; deCesare:2017ynn ; Pithis:2019tvp ; Oriti:2021rvm ; Oriti:2021oux ; Marchetti:2021gcv ; Jercher:2021bie ; Oriti:2023mgu , see also Sections 5.3 and 5.1 of this collection.
@@
which flows from ${d_{\textrm{eff}}}(\xi)=3(r-s)$ at small $\xi$ to ${d_{\textrm{eff}}}(\xi)\to\infty$ at $\xi\to\infty$ (where the $3$ derives from the topological dimension of the $3$-hyperboloid $\mathrm{H}^{3}=\text{SL}(2,\mathbb{C})/\text{SU}(2)$ being the hyperbolic part of $\text{SL}(2,\mathbb{C})$). From this perspective, mean-field theory applies to TGFT on hyperbolic groups because the effective dimension outbalances any finite critical dimension towards large $\xi$, see also Refs. Dekhil:2024ssa ; Dekhil:2024djp for further elaborations. These results provide strong support for the existence of a meaningful continuum gravitational regime in TGFT quantum gravity and the closely related spin foam models as well as lattice quantum gravity and crucially ties it to the presence of the Lorentz group. In particular, the obtained mean-field vacuum open ups the possibility to compare it with the Ashtekar-Lewandowski vacuum Ashtekar:1993wf and the Geiller-Dittrich vacuum Dittrich:2014wpa studied in LQG and spin foam gravity.
@@ 3.4 Observables in Asymptotic Safety
In gravitational theories, measurements are possible due to the presence of additional matter fields which provide a preferred frame of reference DeWitt:1962kpz ; Rovelli:1990pi ; Rovelli:1990ph ; Westman:2007yx ; Goeller:2022rsx . Indeed, the matter fields can be understood as representing the readings on some physical coordinate scaffolding. All measurements of tensor or scalar fields referring to such physical coordinate systems are defined as “relational observables”. Relational observables are therefore natural candidates for composite operators, which can be studied in AS. Moreover at the UV fixed point the corresponding composite operators will have universal scaling exponents. Physically these exponents should appear in the scaling behavior of correlation functions of relational observables at small distances less than the Planck length where effects of the fixed point scaling are expected. As such the computation of these exponents can serve as a way to compare with different approaches to quantum gravity. For example, the same exponents could be computed in other approaches to quantum gravity discussed in this collection, e.g. CDT, GFT or spin foam gravity.
@@ 4 Relational physics in gravity and quantum gravity
Note that the concepts displayed in this section are of particular relevance to the discussions in Section 3 on phase transitions and emergent behavior in quantum gravity, as well as Section 5 on emergent cosmology.
@@ 5 Emergent cosmology from quantum gravity
One might expect such an emergence process to be associated to some form of coarse-graining, and thus to be described via a (non-perturbative) renormalization group analysis of the underlying QG theory Oriti:2013jga ; Carrozza:2016vsq ; Oriti:2021oux ; Marchetti:2022nrf (see also Section 3) but performing such an analysis on a typical QG model is incredibly complicated (it is in fact, already highly non-trivial for most standard QFTs on flat spacetime). On top of that, any physical interpretation of the renormalization procedure is tightly related to spacetime notions. An important example is the renormalization group scale according to which one might want to define the cosmological coarse-graining, and in particular, with respect to which one would describe the average homogeneity of the Universe, see for instance Carrozza:2016vsq ; Marchetti:2020xvf ; Pereira:2019dbn and Eichhorn:2021vid for further discussions.
@@
For such a program to be successful, however, we should be careful to formulate the principles on which our approximations are based in a language appropriate to the fundamental QG theory. Going back to the example of homogeneity, isotropy and cosmic expansion, they should not be formulated in terms of dependence (or independence) with respect to some time (or space) coordinate, as these quantities have completely disappeared in the fundamental QG theory in virtue of background independence. Rather, they should be formulated in relational terms, as dependence (or independence) of physical quantities on various components of a physical reference frame Rovelli:1990ph ; Dittrich:2004cb ; Dittrich:2005kc ; Tambornino:2011vg ; Goeller:2022rsx , see also Section 4 for further details.
All of the above arguments apply in particular to a class of QG models called Group Field Theories (GFTs) Oriti:2009wn ; Oriti:2011jm ; Krajewski:2011zzu ; Freidel:2005qe , see also Sections 3.3 and 5.1 of this collection, and potentially, to a non-perturbative definition of superstring theory, as described by the BFSS matrix model Banks:1996vh . The process of cosmic emergence within these models is reviewed and discussed in Sections 5.3 and 5.4 respectively.
@@ 5.1 On the affinities between Cosmology and Quantum Gravity
Quantum gravity corrections are believed to cure the issue the initial singularity issue that plagues any classical theory of gravity Hawking:1970zqf . Indeed, several quantum gravity proposals can achieve resolution of the initial singularity. One may for instance see this achievement in the context of string theory (e.g., string gas scenario Brandenberger:1988aj (see also Section 5.4) or pre-big-bang cosmology Gasperini:1996fu ) or within a non-perturbative approach to quantum gravity (e.g., loop quantum cosmology bojowald2001absence or group field theory deCesare:2016axk ; deCesare:2017ztf , see also Sections 3.3 and 5.3 of this collection).
@@ 5.3 Emergent Cosmology from (T)GFT Condensates
In $4$ dimensions, the fundamental quanta of GFTs are quantum tetrahedra decorated with discretized geometric and matter field data. Being GFTs proper QFTs, one can construct a Fock space out of the single-tetrahedron Hilbert spaces. The field operator $\hat{\varphi}$ destroying one GFT quantum is then defined on the following domain:
@@
For the reasons mentioned above, performing a non-perturbative renormalization group analysis of GFT models, and thus computing the quantum effective action is currently out of reach. One must therefore look for approximate methods, the simplest one of which being the mean-field approximation Pithis:2018eaq ; Pithis:2019mlv ; Marchetti:2020xvf ; Oriti:2021oux ; Marchetti:2022igl ; Marchetti:2022nrf ; Dekhil:2024ssa ; Dekhil:2024djp . It consists of a saddle point evaluation of the partition function of the system and in the consequent approximation of the effective action via the classical GFT action. From the perspective of the quantum many-body system, this is equivalent to imposing only averaged quantum “equations of motion” on condensate states Gielen:2013naa ; Oriti:2016qtz ; Gielen:2016dss ; Marchetti:2020umh ; Jercher:2021bie , that is
@@
where $\left\langle\cdot\right\rangle_{\sigma}$ represents an average on the state |$\ket{\sigma}$, $\mathop{}\!\mathrm{d}\mathcal{D}$ is an appropriate measure on the domain $\mathcal{D}$ defined above and $\mathcal{N}$ a normalization constant. The above mean-field equations are the QG counterpart of the Gross-Pitaevskii (GP) equation for quantum fluids Oriti:2016qtz As in the theory of quantum fluids, these equations set the stage for a hydrodynamic description of the system pitaevskii2003bose .
@@
It is important to remark, here, that it is at the level of the hydrodynamic variables (and thus of the condensate wavefunction) that one can implement emergent notions as (relational) homogeneity, isotropy, and time evolution. Still, notice that the hydrodynamic variables (8) are in general not relationally localized Marchetti:2020umh .
@@
    The two most important approximations employed are the mean-field approximation and the approximation of negligible interactions. Though in general going beyond mean-field is far from being easy, a promising strategy might be to follow the Bogoliubov approach to quantum fluids pitaevskii2003bose ; david . Including interactions, instead, seems to generally produce cosmic acceleration, which might in turn generate a different, purely quantum geometric mechanism for inflation deCesare:2016rsf ; tom2 , or a phantom dark energy at very late times Oriti:2021rvm .
@@
In conclusion, from the perspective of the fundamental QG theory, relational cosmological physics emerge effectively as a hydrodynamic phenomenon in the limit in which the average number of quanta in the QG system is large. In this sense, a better understanding of the cosmic emergence process goes hand in hand with progresses in condensed matter physics, renormalization, phase transitions and relational physics. On the other hand, QG forces us to revisit these aspects of classical and quantum physics from its own perspective, stripping them of all the redundant, spacetime based information. This offers a unique chance of cross-fertilization between QG and other fields in physics, as discussed in this collection.
@@ 6 Discussion and conclusion
This explained, the collection falls into four thematic units: $(a)$ Map between hydrodynamics and cosmology and analog gravity systems, $(b)$ Phase transitions, continuum limits and emergence in quantum gravity, $(c)$ Relationalism in gravity and quantum gravity and $(d)$ Emergent cosmology from quantum gravity. In the following, we briefly summarize key messages from the individual units:
The first thematic unit, Section 2, reviewed the main ideas and recent results supporting the existence of map between hydrodynamics and cosmology on the one hand, and the possibility to realize analog systems for quantum cosmology in the future on the other. The first contribution discussed the shared symmetries between hydrodynamics and symmetry-reduced gravitational superspaces. Crucially, these shared symmetries allow to build a dictionary between the observables of quantum cosmology and the ones of hydrodynamics opening the door for analog models to encode dynamical information. The second contribution summarized the efforts to consider the backreaction of quantum fluctuations on the background in analog gravity. The third contribution finally reviewed the status of analog models tailored to cosmological settings. Importantly, preliminary results on an analog BEC model for reproducing the WDW dynamics show that the geometry of the field space is indeed close to a cosmological spacetime geometry.
@@
The fourth thematic unit, Section 5, examined the process of cosmic emergence in quantum gravity, highlighting the challenges inherent in reconstructing continuum cosmological concepts from a pre-geometric quantum gravity framework, as well as the significant impact quantum gravity can have on cosmology by helping to alleviate both conceptual and observational tensions. The first contribution focused on this potential impact, with particular attention to the complex interplay between quantum gravity, inflation, and dark energy. In contrast, the second contribution addressed the challenges associated with emergence itself, specifically the subtle issues surrounding the emergence of time. Finally, concrete examples of cosmic emergence were provided in the third and fourth contributions, within the TGFT approach and superstring theory, respectively. These examples demonstrated how macroscopic properties of the universe can be quantitatively linked to microscopic quantum-gravitational phenomena.

### v2

@@ 3.1 Effective dynamics for the spatial volume of 3-dimensional CDT
One feature of CDT is the existence of a phase characterized by the emergence of a macroscopic universe, suggestive of a semiclassical behavior at large scales. A spacetime condensation phenomenon underlies the emergence of such a semiclassical universe, where the time extension of the condensate, or droplet, is strictly smaller than the total time of the quantum universe. In Benedetti:2014dra ; Benedetti:2016rwo ; Benedetti:2022ots we introduced an effective model for the dynamics of the spatial volume in $2+1$ dimensions, which can be obtained as a mini-superspace reduction of Hořava-Lifshitz gravity, and we have shown that a study of the minima of the action (supported also by Monte-Carlo simulations) leads to a phase diagram with a droplet phase similar to that of CDT, which instead cannot be reproduced with a model based on a mini-superspace reduction of Einstein’s gravity theory. Therefore, in view also of previous results  Horava:2009if ; Benedetti:2009ge ; Ambjorn:2010hu ; Budd:2011zm ; Ambjorn:2013joa , we conclude that there is strong evidence that the effective field theory associated to the continuum limit of CDT in $1+1$ and $2+1$ dimensions is a foliation-preserving gravity theory, that is, a theory in the same general class to which Hořava-Lifshitz gravity belongs (see Benedetti:2022ots for a detailed argument, where the effective mini-superspace models are interpreted as a mean-field theory of CDT).
@@ 3.2 Spin foams, refinement limit and semi-classical physics
On a more speculative note, the question arises how cosmology may emerge from the full theory before imposing symmetry restrictions. Cosmological dynamics may emerge (on average) by considering coarse-grained observables, like the 3-volume of a spatial slice, and matter fields as relational clocks and rods, revealing similarities and complementary perspectives to the hydrodynamics on superspace program.
@@ 3.3 Three Ways to Mean-Field Group Field Theory
Mean-field theory is applicable if ${d_{\textrm{eff}}}>{d_{\textrm{crit}}}$. This works even for compact group if there are sufficiently many matter degrees of freedom adding to ${d_{\textrm{eff}}}$. On groups with hyperbolic geometry, as natural when there is some Lorentzian structure in a TGFT model as in Jercher:2021bie ; Jercher:2022mky , mean-field theory applies generically as ${d_{\textrm{eff}}}(k)\to\infty$ excedes the critical dimension in any case Marchetti:2022igl ; Marchetti:2022nrf . This lays the theoretical foundations for applications of TGFT to phenomenology based on condensate states in a mean-field regime Gielen:2013kla ; Oriti:2015rwa . In particular, these results strongly lend support to the TGFT condensate cosmology program which has successfully extracted effective continuum cosmological dynamics from TGFT condensate states Gielen:2013kla ; Gielen:2013naa ; deCesare:2016rsf ; Pithis:2016wzf ; Pithis:2016cxg ; deCesare:2017ynn ; Pithis:2019tvp ; Oriti:2021rvm ; Oriti:2021oux ; Marchetti:2021gcv ; Jercher:2021bie ; Oriti:2023mgu , see also Sections 5.1 and 5.3 of this collection. As explained in the latter as well as in the introduction to this collection, this program is a particular instantiation of the hydrodynamics on superspace framework.
@@
which flows from ${d_{\textrm{eff}}}(\xi)=3(r-s)$ at small $\xi$ to ${d_{\textrm{eff}}}(\xi)\to\infty$ at $\xi\to\infty$ (where the $3$ derives from the topological dimension of the $3$-hyperboloid $\mathrm{H}^{3}=\text{SL}(2,\mathbb{C})/\text{SU}(2)$ being the hyperbolic part of $\text{SL}(2,\mathbb{C})$). From this perspective, mean-field theory applies to TGFT on hyperbolic groups because the effective dimension outbalances any finite critical dimension towards large $\xi$, see also Refs. Dekhil:2024ssa ; Dekhil:2024djp for further elaborations. These results provide strong support for the existence of a meaningful continuum gravitational regime in TGFT quantum gravity and the closely related spin foam models as well as lattice quantum gravity and crucially ties it to the presence of the Lorentz group. In particular, the obtained mean-field vacuum opens up the possibility to compare it with the Ashtekar-Lewandowski vacuum Ashtekar:1993wf and the Geiller-Dittrich vacuum Dittrich:2014wpa studied in LQG and spin foam gravity.
@@ 3.4 Observables in Asymptotic Safety
In gravitational theories, measurements are possible due to the presence of additional matter fields which provide a preferred frame of reference DeWitt:1962kpz ; Rovelli:1990pi ; Rovelli:1990ph ; Westman:2007yx ; Goeller:2022rsx , see in particular the contributions in Section 4 on relationalism. Indeed, the matter fields can be understood as representing the readings on some physical coordinate scaffolding. All measurements of tensor or scalar fields referring to such physical coordinate systems are defined as “relational observables”. Relational observables are therefore natural candidates for composite operators, which can be studied in AS. Moreover at the UV fixed point the corresponding composite operators will have universal scaling exponents. Physically these exponents should appear in the scaling behavior of correlation functions of relational observables at small distances less than the Planck length where effects of the fixed point scaling are expected. As such the computation of these exponents can serve as a way to compare with different approaches to quantum gravity. For example, the same exponents could be computed in other approaches to quantum gravity discussed in this collection, e.g. CDT, GFT or spin foam gravity.
@@ 4 Relational physics in gravity and quantum gravity
Note that while the contributions below can be viewed as standalone perspectives on their respective topics, they collectively contribute to the development of the relational strategy. In particular, they advocate for a notion of physical localization in field space rather than spacetime, which is an essential component of the hydrodynamics-on-superspace framework. In this way, the contributions also feed into its further development. Finally, note that the concepts displayed in this section are of particular relevance to the discussions in Section 3 on phase transitions and emergent behavior in quantum gravity, as well as Section 5 on emergent cosmology (see in particular Section 5.3).
@@ 5 Emergent cosmology from quantum gravity
One might expect such an emergence process to be associated to some form of coarse-graining, and thus to be described via mean-field or (non-perturbative) renormalization group analyses of the underlying QG theory Oriti:2013jga ; Carrozza:2016vsq ; Oriti:2021oux ; Marchetti:2022nrf ; Eichhorn:2018phj ; Benedetti:2022ots (see also Section 3) but performing such an analysis on a typical QG model is incredibly complicated (it is in fact, already highly non-trivial for most standard QFTs on flat spacetime). On top of that, any physical interpretation of the coarse-graining procedure is tightly related to spacetime notions. An important example is the renormalization group scale according to which one might want to define the cosmological coarse-graining, and in particular, with respect to which one would describe the average homogeneity of the Universe, see for instance Carrozza:2016vsq ; Marchetti:2020xvf ; Pereira:2019dbn and Eichhorn:2021vid for further discussions.
@@
For such a program to be successful, however, we should be careful to formulate the principles on which our approximations are based in a language appropriate to the fundamental QG theory. Going back to the example of homogeneity, isotropy and cosmic expansion, they should not be formulated in terms of dependence (or independence) with respect to some time (or space) coordinate, as these quantities have completely disappeared in the fundamental QG theory in virtue of background independence. Rather, they should be formulated in relational terms, as dependence (or independence) of physical quantities on various components of a physical reference frame Rovelli:1990ph ; Dittrich:2004cb ; Dittrich:2005kc ; Tambornino:2011vg ; Goeller:2022rsx , see also Section 4 for further details on the relationalist perspective.
All of the above arguments apply in particular to a class of QG models called Group Field Theories (GFTs)  Oriti:2009wn ; Oriti:2011jm ; Krajewski:2011zzu ; Freidel:2005qe . The process of cosmic emergence within this approach naturally fits into the hydrodynamics on superspace framework, and is reviewed and discussed in Section 5.3. Some key related results are also highlighted in Section 5.1. These findings are directly connected to the contribution on condensate states in GFT presented in Section 3.3 of this collection.
Potentially, the arguments outlined in this introduction also apply to a non-perturbative definition of superstring theory, as described by the BFSS matrix model Banks:1996vh which is discussed in a cosmological context in Section 5.4. It is left to future research to investigate how the latter section could fit in detail into this novel perspective on cosmology from quantum gravity.
@@ 5.1 On the affinities between Cosmology and Quantum Gravity
Quantum gravity corrections are believed to cure the issue of the initial singularity that plagues any classical theory of gravity Hawking:1970zqf . Indeed, several quantum gravity proposals can achieve resolution of the initial singularity. One may for instance see this achievement in the context of string theory (e.g., string gas scenario Brandenberger:1988aj (see also Section 5.4) or pre-big-bang cosmology Gasperini:1996fu ) or within a non-perturbative approach to quantum gravity (e.g., loop quantum cosmology bojowald2001absence or group field theory deCesare:2016axk ; deCesare:2017ztf , see also Sections 3.3 and 5.3 of this collection).
@@ 5.3 Emergent Cosmology from (T)GFT Condensates
The following contribution gives a lightning review of the TGFT condensate cosmology approach. This will clarify how the topic of emergence of continuum geometry is addressed therein, see also Section 3.3, and how the relationalist perspective, discussed in Section 4, is exploited to provide an approximate description thereof together with a deparametrized account of its dynamics in a cosmological context. In this way, it explains how this approach corresponds to a specific realization of the hydrodynamics on superspace framework.
To start off, consider that in $4$ dimensions the fundamental quanta of GFTs are quantum tetrahedra decorated with discretized geometric and matter field data. Being GFTs proper QFTs, one can construct a Fock space out of the single-tetrahedron Hilbert spaces Oriti:2013aqa ; Gielen:2013naa ; Jercher:2021bie ; Sahlmann:2023plc . The field operator $\hat{\varphi}$ destroying one GFT quantum is then defined on the following domain:
@@
For the reasons mentioned above, performing a non-perturbative renormalization group analysis of GFT models, and thus computing the quantum effective action is currently out of reach. One must therefore look for approximate methods, the simplest one of which being the mean-field approximation [^4] We refer the reader to Section 3.3 for a discussion on the validity of the mean-field approximation.  Pithis:2018eaq ; Pithis:2019mlv ; Marchetti:2020xvf ; Oriti:2021oux ; Marchetti:2022igl ; Marchetti:2022nrf ; Dekhil:2024ssa ; Dekhil:2024djp . It consists of a saddle point evaluation of the partition function of the system and in the consequent approximation of the effective action via the classical GFT action. From the perspective of the quantum many-body system, this is equivalent to imposing only averaged quantum “equations of motion” on condensate states Gielen:2013naa ; Oriti:2016qtz ; Gielen:2016dss ; Marchetti:2020umh ; Jercher:2021bie , that is
@@
where $\left\langle\cdot\right\rangle_{\sigma}$ represents an average on the state |$\ket{\sigma}$, $\mathop{}\!\mathrm{d}\mathcal{D}$ is an appropriate measure on the domain $\mathcal{D}$ defined above and $\mathcal{N}$ a normalization constant. The above mean-field equations are the QG counterpart of the Gross-Pitaevskii (GP) equation for quantum fluids Oriti:2016qtz . As in the theory of quantum fluids, these equations set the stage for a hydrodynamic description of the system pitaevskii2003bose .
@@
It is important to remark, here, that it is at the level of the hydrodynamic variables (and thus of the condensate wavefunction) that one can implement emergent notions as (relational) homogeneity, isotropy, and time evolution. Still, notice that the hydrodynamic variables (8) are in general not relationally localized Marchetti:2020umh ; lucaed .
@@
    At the level of the quantum dynamics, the two most important approximations employed are the mean-field approximation and the approximation of negligible interactions. Though in general going beyond mean-field is far from being easy, a promising strategy might be to follow the Bogoliubov approach to quantum fluids pitaevskii2003bose ; david . Including interactions, instead, seems to generally produce cosmic acceleration, which might in turn generate a different, purely quantum geometric mechanism for inflation deCesare:2016rsf ; Pithis:2016cxg ; Jercher:2021bie ; tom2 , or a phantom dark energy at very late times Oriti:2021rvm .
    Regarding the implementation of the relational strategy, exact relational observables have been constructed for the first time in the GFT Fock space by incorporating insights gained within the quantum reference frame context (see Section 4) into the GFT framework lucaed . These relational observables, constructed using frame POVMs, match the effective relational observables described above when the frame measurement associated with the POVMs carries intrinsic uncertainty lucaed .
@@
In conclusion, from the perspective of the fundamental QG theory, relational cosmological physics effectively emerge as a hydrodynamic phenomenon in the limit in which the average number of quanta in the QG system is large. This section shows that the novel hydrodynamics on superspace framework, as advertised by the editors of this collection in the introduction and Section 2.1, can be concretely realized for particular TGFT models via a mean-field approximation of the fundamental quantum gravity dynamics (see Section 3.3) and a relational description of the emergent physics (see Section 4). In this way, the work presented in Section 5.3 may also serve as a template for extracting cosmological physics from superspace hydrodynamics (including QG effects) from the point of view of other QG approaches.
@@ 6 Discussion and conclusion
This explained, the collection falls into four thematic units: $(a)$ Map between hydrodynamics and cosmology and analog gravity systems, $(b)$ Phase transitions, continuum limits and emergence in quantum gravity, $(c)$ Relationalism in gravity and quantum gravity and $(d)$ Emergent cosmology from quantum gravity. The content of these units may be seen as standalone perspectives on their respective topics. Beyond this, the results presented therein inform the development of the hydrodynamics on superspace framework. In the following, we briefly summarize key messages from the individual units:
The first thematic unit, Section 2, reviewed the main ideas and recent results supporting the existence of a map between hydrodynamics and cosmology on the one hand, and the possibility to realize analog systems for quantum cosmology in the future on the other. The first contribution discussed the shared symmetries between hydrodynamics and symmetry-reduced gravitational superspaces. Crucially, these shared symmetries allow to build a dictionary between the observables of quantum cosmology and the ones of hydrodynamics opening the door for analog models to encode dynamical information. The second contribution summarized the efforts to consider the backreaction of quantum fluctuations on the background in analog gravity. In a sense, this effort goes in the same direction as it aims at encoding effects of the dynamics on the background geometry. The third contribution finally reviewed the status of analog models tailored to cosmological settings. Importantly, preliminary results on an analog BEC model for reproducing the WDW dynamics show that the geometry of the field space is indeed close to a cosmological spacetime geometry.
@@
The fourth thematic unit, Section 5, examined the process of cosmic emergence in quantum gravity, highlighting the challenges inherent in reconstructing continuum cosmological concepts from a pre-geometric quantum gravity framework, as well as the significant impact quantum gravity can have on cosmology by helping to alleviate both conceptual and observational tensions. The first contribution focused on this potential impact, with particular attention to the complex interplay between quantum gravity, inflation, and dark energy. In contrast, the second contribution addressed the challenges associated with emergence itself, specifically the subtle issues surrounding the emergence of time. Finally, concrete examples of cosmic emergence were provided in the third and fourth contributions, within the TGFT approach and superstring theory, respectively. These examples demonstrated how macroscopic properties of the universe can be quantitatively linked to microscopic quantum-gravitational phenomena. In particular, this unit demonstrated that TGFT condensate cosmology is a specific instantiation of the hydrodynamics on superspace framework, as advertised by the editors of this collection, and may serve as a template for the extraction of cosmology from superspace hydrodynamics, which encompasses QG signatures, from other quantum gravity approaches.